OpenStudy (anonymous):

Will fan and medal! sec(-pi/3) - cot(-5pi/4) I got -3 instead of 3 Cant use calculators

OpenStudy (nathanjhw):

Yes 3 is correct.

OpenStudy (anonymous):

I dont know where I missed up when I did my math cause I got -3

OpenStudy (anonymous):

@NathanJHW

OpenStudy (nathanjhw):

You most likely just missed a negative sign.

OpenStudy (nathanjhw):

Try it again from scratch.

OpenStudy (anonymous):

At the last bit, is it supposed to be -2-1, when you're doing the work

OpenStudy (anonymous):

@NathanJHW

OpenStudy (nathanjhw):

Yes.

OpenStudy (nathanjhw):

No sorry. It's 2 -(-1)

OpenStudy (nathanjhw):

sec(-pi/3) = 2 cot(-5pi/4) = -1

OpenStudy (anonymous):

Do the negatives in the original equation matter?

OpenStudy (nathanjhw):

Yes because 2 - (-1) = 3 and 2 -1 = 1

OpenStudy (dtan5457):

@minisweet4 Look at sec(-pi/3) I'd advise you to solve for the inverse, cos first. Notice this is NEGATIVE 60 degrees...That is the 4th quadrant...however in the 4th quadrant, Cos is POSITIVE. You do your 30-60-90 triangle (i'm assuming you know this already) and you get 1/2 as cos, so sec is 2. NEXT cot(-5pi/4) That is negative 225 degrees, which ends up in the 2nd quadrant, NEGATIVE for cot and tan. do your 45-45-90 triangle (again i'm assuming you know this, feel free to ask if you don't) and you get 1, obviously -1 since this quadrant is negative. 2-(-)1=3

OpenStudy (dtan5457):

So basically when you have negative degrees, that doesn't necessarily make it negative, it all depends on the quadrant you end up in.